OpenStudy (anonymous):

Two numbers are w and 3w - 5.  Their sum is 23.  Solve the equation w + 3w -5 = 23 to find the numbers.

OpenStudy (anonymous):

4w-5=23 4w=23+5 4w=28 w=28/4 w=7